Suppose that a small national economy consists of one firm. During a certain period of time the firm: 

Imports of raw material from abroad costing   Rs. 4000

Hires labor, who are paid wages are Rs. 9000

Sells all its output for Rs. 20000 and so makes a profit of Rs. 7000

Pays its post-tax profit of Rs. 4000 to stockholders 




The country‘s government taxes the labor force Rs. 2000 and the company Rs. 3000

The firm‘s sales of Rs. 20000 are accounted for by:

Domestic consumers who spend Rs.11000 this Rs.11000 is the post-tax wages earned by the labor force (Rs.7000) plus Rs.4000 in dividends earned by the company company’s stockholders;

The government which spends the Rs. 5000 it has raised in taxes ;

Foreign buyers who spend Rs. 4000

Calculate the GDP with three measures

Income Approach

Expenditure Approach

Output or value-added Approach


1
Expert's answer
2021-04-20T17:25:17-0400

Income Approach

Y = Wages + Profits = 9,000 + 7,000 = 16,000.

Expenditure Approach

Y = C + I + G + NX = 11,000 + 4,000 + 5,000 - 4000 = 16,000.

Output or value-added Approach

Y = Gross value of output - Value of intermediate consumption = 20,000 - 4,000 = 16,000.